How many grams of potassium phosphate will be formed upon the complete reaction of 30.7 grams of phosphoric acid with excess potassium hydroxide
1
Expert's answer
2017-03-06T13:34:05-0500
H₃PO₄+3KOH=K₃PO₄+3H₂O

1MolH₃PO₄ -> 1Mol K₃PO₄

V(H₃PO₄)=m/M=30.7g/98g/Mol=0.313Mol

V(K₃PO₄)=0.313Mol

m(K₃PO₄)=V*M=0.313Mol*212g/Mol=66.36g